Circular No. 8630

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2005kd, 2005ke, 2005kf Three apparent supernovae have been reported: 2005kd by T. Puckett

and A. Pelloni (via the 0.35-m automated supernova patrol telescope; cf. IAUC 8629), and 2005ke and 2005kf by M. Baek, R. R. Prasad, and W. Li (via LOSS/KAIT; cf. IAUC 8628).

SN  2005 UT  α_{2000}  δ_{2000}  Mag.  Offset
2005kd Nov. 12.22 4^{h}03^{m}16.^{s}88 +71^{o}43′18′′.9 17.0 0′′.1 W, 5′′.0 N
2005ke Nov. 13.33 3 35 04.35 −24 56 38.8 17.2 40′′ E, 40′′ S
2005kf Nov. 11.53 7 47 26.51 +26 55 32.4 17.2 0′′.8 E, 0′′.6 S

Additional approximate magnitudes for 2005kd in PGC 14370: Sept. 18 and Nov. 9, [20.0; Nov. 13.02 UT, 16.7. Additional approximate magnitudes for 2005ke in NGC 1371: Oct. 24.32, [19.5; Nov. 14.35, 16.3. Additional approximate magnitudes for 2005kf, which is in a galaxy located 82′′ east and 16′′.5 south of the center of NGC 2449: Apr. 5.20, [19.5; Oct. 14.52, 17.0; 20.53, 16.5; 31.50, 16.7; Nov. 16.52, 17.5.

SN 2005kd has been found by an Ohio State University group (J. Eastman et al., communicated by J. Prieto; cf. CBET 290) to be a young type-IIn supernova from a spectrogram (range 390–730 nm) taken on Nov. 13.3 UT with the MDM 2.4-m telescope (+ CCDS); the spectrum shows a blue continuum and strong hydrogen Balmer (Hα, Hβ, Hγ, and Hδ) and He I lines in emission at central wavelengths fully consistent with the recession velocity of PGC 14370.

(2006) POLONSKAYA D. Pray, Greene, RI; P. Pravec and P. Kušnirák, Ondřejov Observatory;

and W. Cooney, J. Gross and D. Terrell, Sonoita Research Observatory, report that observations obtained during Nov. 1–13 show that the minor planet (2006) consists of a pair of bodies orbiting each other with a period of 19.15 ± 0.02 hr and producing mutual eclipse/occultation events 0.06-mag deep. There are present two rotational lightcurves with periods of 3.1183 ± 0.0002 and 6.656 ± 0.001 hr, with amplitudes of 0.08 and 0.06 mag, respectively, that persist during the mutual events. This suggests that either the system is binary with both bodies rotating non-synchronously with the orbital motion or there is a third body producing the additional period. A lower limit on the mean-diameter ratio of the eclipsing pair is 0.22. Further observations are needed to fully resolve the system.

2005 November 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green